OpenStudy (anonymous):

how many moles is 1.204 x 10^24?

OpenStudy (anonymous):

\[?mol = 1.204\times 10^{24} \times \frac{ 1 mol }{ 6.022 \times 10^{23} } \approx 2 mol\]